The Sports Authority, Inc. is a private full-line sporting goods retailer. Assume one of the Sports Authority stores reported current assets of $88,550 and its current ratio was 1.75, and then completed the following transactions: (1) paid $5,700 on accounts payable, (2) purchased a delivery truck for $11,500 cash, (3) wrote off a bad account receivable for $2,300, and (4) paid previously declared dividends in the amount of $26,500. Required: Compute the updated current ratio after each transaction, by showing the cumulative effects of the transactions in the following table. (Round your answers to 2 decimal places.) Current Ratio 1.85 Transaction (1) Transaction (2) Transaction (3) Transaction (4) 1.59 Hints ReferenceseBook & Resources Hint #1
Answer #1
Ans. Current Ratio = CA / CL
1.75 = 88,550 / CL
CL = 88,550 / 1.75
= $             50,600
Transaction (1)
Current Ratio = CA / CL
= (88,550-5,700)/ (50,600-5,700)
= 82,850 / 44,900
= 1.85
Transaction (2)
Current Ratio = CA / CL
= (88,550 -11,500)/ 50,600
= 77,050 / 50,600
= 1.52
Transaction (3)
Current Ratio = CA / CL
= (88,550 -2,300)/ 50,600
= 86,250 / 50,600
= 1.70
Transaction (4)
Current Ratio = CA / CL
= (88,550 -26,500)/ 50,600
= 62,050 / 50,600
= 1.23